Consider a 15,000-Ci 60Co source which gives off gamma rays of 1.17 MeV and 1.33 MeV from each decay. Assume that the dose rate at a distance R (cm) from a source of strength C (Curies) emitting gamma energy E (MeV) for each disintegration is given by:

D(mrad/h) = 4.6 x 10^6 CE/R^2

Calculate the dose rate at 1 m from the source and the time to receive the 75 rad. (Ans. t=21 seconds)
Answer #1

Strength of the source, C = 15,000 Ci

Total energy of Gamma rays emitted from one decay, E=1.17 +1.33 MeV =2.50 MeV

Distance R = 1m = 100 cm

Dose rate

D=4.6*10^6 \frac{CE}{R^2}=4.6*10^6 *\frac{15000*2.50}{100^2} mrad/h

D=1.72*10^7 mrad/h

D=\frac{1.72*10^7*10^{-3}}{3600} rad/s

D=4.8 rad/s

Time to receive a dose of Dt=75 rad is:

t=\frac{D_t}{D}=\frac{75}{4.8}s = 15.6s